# Craw4LLM: Efficient Web Crawling for LLM Pretraining

### v2

@@ Abstract
Shi Yu12thanks: Work done while visiting Carnegie Mellon University.  Zhiyuan Liu1  Chenyan Xiong2 1Department of Computer Science and Technology, Tsinghua University 2School of Computer Science, Carnegie Mellon University yus21@mails.tsinghua.edu.cn; liuzy@tsinghua.edu.cn; cx@cs.cmu.edu
@@ 1 Introduction
We conduct large-scale crawling simulations on ClueWeb22-A (Overwijk et al., 2022), a snapshot of the web containing 900 million English webpages obtained from the central index of a commercial search engine. Results show that, by crawling only 1× of the pretraining dataset size, Craw4LLM can outperform traditional crawlers which collect 1×, 2×, and 4× more data followed by data selection. Compared to the baseline crawler that achieves the same performance, Craw4LLM crawls only 21% of the webpages. Further analysis reveals that during crawling, Craw4LLM quickly discovers documents that align with the oracle selection, which crawls the full web graph. As a result, it achieves 95% of the oracle performance while crawling only 2.2% of the data.
@@ 2 Methodology
Commonsense Reasoning Language Understanding Reading Comprehension Symbolic Problem Solving World Knowledge Core % of
Crawling Method Selection Pool Size (4 tasks) (6 tasks) (3 tasks) (5 tasks) (5 tasks) (23 tasks) Oracle
Oracle Selection (Upper Bound): Random sample from the top 10% rated data from ClueWeb22 using DCLM fastText for pretraining
n.a. 45× 0.2438 0.2209 0.1483 0.2039 0.2403 0.2239 100%
Crawl-then-Select: Crawl 1× and 2× more data from ClueWeb22 and select top-rated 1× data using DCLM fastText for pretraining
Random 0.1906 0.1890 0.0244 0.1834 0.1930 0.1748 78.1%
@@
Ours: Crawl 1× data using Craw4LLM for pretraining
Craw4LLM 0.2116 0.2311 0.0826 0.1979 0.2486 0.2133 95.3%
Table 1: Downstream LLM performance. All models are pretrained on 1× data, which corresponds to 20M documents and 32.9B tokens. The evaluation metric is centered accuracy (0 = random guess) (Li et al., 2024). Best/2nd best in the last two groups are bolded/underlined. See Appendix C for detailed results.
@@ 3 Experimental Methodology
To run experiments in our limited computational budget, we run a simulation of Craw4LLM on the ClueWeb22 dataset (Overwijk et al., 2022), a snapshot of the web with graph information from a commercial crawler. We use the English subset of ClueWeb22-A, which is a web graph containing 900M webpages with links. We randomly sampled 10K URLs as our seed URLs. We set the number of total crawled documents $N$ to 20M and crawled documents each iteration $n$ to 10K. We use the DCLM fastText classifier (Li et al., 2024) as the pretraining influence scorer $\mathcal{M}(\cdot)$.
@@
We emulate traditional graph-connectivity-based crawlers by replacing the LLM-oriented URL scoring function (Eq. 1) with a function that returns the indegree for a given URL, since a node’s indegree closely correlates with PageRank, a common graph connectivity metric, as shown in Figure 2(b) and previous findings (Fortunato et al., 2008). We also introduce a random crawling baseline, where the scorer assigns random scores. We run both of them in a crawl-then-select setting, first crawling 1× or 2× more documents and then selecting the top 1× (20M) documents based on scores assigned by the DCLM fastText classifier. This process mimics existing data-filtering pipelines, which begin with crawled documents and then apply filtering (Li et al., 2024; Penedo et al., 2024).
@@
We also introduce an oracle selection run in which we directly apply the DCLM fastText classifier to the entire ClueWeb22-A document set and select the top 10% documents for pretraining, serving as the upper bound.
@@ 4 Evaluation Results
In this section, we first present the overall performance of Craw4LLM (Sec. 4.1), followed by further analysis (Sec. 4.2).
@@ 4.1 Overall Performance
In this experiment, we compare the performance of Craw4LLM with baseline crawlers by evaluating LLMs trained on their respective crawled data. As shown in Table 1, when all methods crawl the same amount of training data (1×), Craw4LLM significantly outperforms random crawling and indegree crawling. In the crawl-then-select setting, where traditional crawlers are allowed to collect twice as much data (2×) for later selection, they still underperform compared to Craw4LLM. This suggests that incorporating pretraining-oriented signals early in the crawling process is more beneficial than relying on post-selection. With only 1× of the data, Craw4LLM retains 95% of the performance achieved by the oracle run, which directly selects from a substantially larger 45× data pool.
In Section 4.2, we further analyze the efficiency of Craw4LLM compared to traditional crawlers and explore the reasons behind it.
@@ 4.2 Analysis
We evaluate the efficiency of Craw4LLM by comparing the number of documents it crawls or visits against baseline crawlers. As shown in Figure 3(a), even when the baselines crawl 4× the required pretraining data for selection, they still underperform compared to Craw4LLM. Extrapolation suggests that the indegree-based crawler would need to process 4.8× more documents (96M) to match Craw4LLM’s performance. Figure 3(b) further illustrates that Craw4LLM achieves the same performance while crawling only 21% of the documents required by the indegree-based crawler, or 48% when considering all visited documents. These results highlight the efficiency of Craw4LLM, demonstrating its potential to reduce website burdens and mitigate over-crawling.
@@ Appendix A Details on Crawling
We run our simulated crawlers on a Linux server equipped with two Intel(R) Xeon(R) E5-2630 v3 CPUs (8 cores per socket, 16 cores in total, 1 thread per core), 125GiB of memory, and an SSD. A crawl of 20 million documents takes approximately one day to complete.
@@ Appendix B Details on LLM Training and Evaluation
Appendix C Detailed Results
Crawling Selection Commonsense Reasoning
Method Pool Size CommonsenseQA COPA OpenBookQA PIQA
Oracle Selection (Upper Bound)
n.a. 45× 0.2850 0.7000 0.3300 0.6812
Crawl-then-Select
@@ Appendix C Detailed Results
Random 0.2588 0.6200 0.3160 0.6785
Random 0.2326 0.6400 0.3380 0.6757
@@
Indegree 0.1966 0.6600 0.3040 0.6752
Indegree 0.2088 0.6400 0.3400 0.6817
Ours
Craw4LLM 0.2277 0.6600 0.3300 0.6926
@@
Crawling Selection Language Understanding
Method Pool Size BIG-Bench Lang. Id. HellaSwag (zero-shot) HellaSwag LAMBADA Winograd Winogrande
Oracle Selection (Upper Bound)
n.a. 45× 0.2515 0.3856 0.3905 0.4432 0.6557 0.5130
Crawl-then-Select
@@
Random 0.2468 0.3882 0.3925 0.4073 0.6007 0.5130
Random 0.2521 0.4011 0.4019 0.4390 0.6154 0.5130
@@
Indegree 0.2547 0.3749 0.3771 0.3773 0.5861 0.5241
Indegree 0.2562 0.3994 0.4008 0.4159 0.6190 0.5178
Ours
Craw4LLM 0.2544 0.4035 0.4048 0.4196 0.6593 0.5288
@@
Crawling Selection Reading Comprehension
Method Pool Size BoolQ CoQA SQuAD
Oracle Selection (Upper Bound)
n.a. 45× 0.5755 0.2479 0.3139
Crawl-then-Select
@@
Random 0.5807 0.2053 0.2759
Random 0.5911 0.2361 0.2951
@@
Indegree 0.5697 0.1843 0.2390
Indegree 0.5765 0.2147 0.2736
Ours
Craw4LLM 0.5440 0.2264 0.2215
@@
Crawling Selection Symbolic Problem Solving
Method Pool Size AGI Eval LSAT-AR BIG-Bench CS Algorithms BIG-Bench Dyck Lang. BIG-Bench Operators BIG-Bench Repeat Copy Logic
Oracle Selection (Upper Bound)
n.a. 45× 0.2739 0.4341 0.2160 0.2143 0.0625
Crawl-then-Select
@@
Random 0.2696 0.4538 0.2520 0.1762 0.0313
Random 0.1957 0.4568 0.2600 0.1857 0.0625
@@
Indegree 0.2609 0.4235 0.2340 0.2143 0.0313
Indegree 0.2174 0.4538 0.2530 0.1667 0.0938
Ours
Craw4LLM 0.2696 0.4371 0.1620 0.2095 0.0938
@@
Crawling Selection World Knowledge
Method Pool Size ARC Easy ARC Challenge BIG-Bench QA Wikidata Jeopardy MMLU
Oracle Selection (Upper Bound)
n.a. 45× 0.5951 0.3166 0.4945 0.1176 0.2805
Crawl-then-Select
@@
Random 0.5425 0.2807 0.5081 0.0648 0.2561
Random 0.5577 0.2867 0.5126 0.0970 0.2543
@@
Indegree 0.5248 0.2790 0.5205 0.0555 0.2464
Indegree 0.5749 0.2935 0.5084 0.0959 0.2430
Ours
Craw4LLM 0.6103 0.3208 0.5143 0.1323 0.2661
@@
Appendix D The ClueWeb22 Dataset

### v3

@@ Abstract
Shi Yu12thanks: Work done while visiting Carnegie Mellon University.  Zhiyuan Liu1  Chenyan Xiong2 1Department of Computer Science and Technology, Tsinghua University 2Language Technologies Institute, Carnegie Mellon University yus21@mails.tsinghua.edu.cn; liuzy@tsinghua.edu.cn; cx@cs.cmu.edu
@@ 1 Introduction
We conduct large-scale crawling simulations on ClueWeb22-A (Overwijk et al., 2022), a snapshot of the web containing 900 million English webpages obtained from the central index of a commercial search engine. Results show that, by crawling only 1× of the pretraining dataset size, Craw4LLM can outperform traditional crawlers which collect 1×, 2×, and 4× data followed by data selection. Compared to the baseline crawler that achieves the same performance, Craw4LLM crawls only 21% of the webpages. Further analysis reveals that during crawling, Craw4LLM quickly discovers documents that align with the oracle selection, which selects from the full web graph. As a result, it achieves over 95% of the oracle performance while crawling only 2.2% of the data.
@@ 2 Methodology
Selection Commonsense Reasoning Language Understanding Reading Comprehension Symbolic Problem Solving World Knowledge Core % of
Method Pool Size (4 tasks) (6 tasks) (3 tasks) (5 tasks) (5 tasks) (23 tasks) Oracle
Using the DCLM fastText (Li et al., 2024) classifier as the pretraining influence scorer
Oracle 45× 0.2438 0.2209 0.1483 0.2039 0.2403 0.2239 100%
[…]
@@
[…]
Using the FineWeb-Edu (Penedo et al., 2024) classifier as the pretraining influence scorer
Oracle 45× 0.1899 0.1973 0.1081 0.2117 0.2786 0.2133 100%
[…]
0.1797 0.1888 0.0931 0.1586 0.2100 0.1807 84.7%
Indegree 0.1730 0.1680 0.0326 0.1616 0.1668 0.1556 72.9%
0.1720 0.1709 0.0840 0.1783 0.1842 0.1724 80.8%
[…]
Table 1: Downstream LLM performance. Either the DCLM fastText classifier (Li et al., 2024) (top) or FineWeb-Edu classifier (Penedo et al., 2024) (down) is used as the pretraining influence scorer for Craw4LLM, and to select documents for oracle and crawl-then-select runs (Section 3). All models are pretrained on 1× data (20M documents, 32.9B tokens). The evaluation metric is centered accuracy (0 = random guess) (Li et al., 2024). Best/2nd best in the last two groups are bolded/underlined. See Appendix D for detailed results.
@@ 3 Experimental Methodology
To run experiments in our limited computational budget, we run a simulation of Craw4LLM on the ClueWeb22 dataset (Overwijk et al., 2022), a snapshot of the web with graph information from a commercial crawler. We use the English subset of ClueWeb22-A, which is a web graph containing 900M webpages with links. We randomly sampled 10K URLs as our seed URLs. We set the number of total crawled documents $N$ to 20M and crawled documents each iteration $n$ to 10K. In our experiments, we consider using the DCLM fastText classifier (Li et al., 2024) or the FineWeb-Edu classifier (Penedo et al., 2024) as the pretraining influence scorer $\mathcal{M}(\cdot)$, whose details can be found in Appendix C.
@@
We emulate traditional graph-connectivity-based crawlers by replacing the LLM-oriented URL scoring function (Eq. 1) with a function that returns the indegree for a given URL, since a node’s indegree closely correlates with PageRank, a common graph connectivity metric, as shown in Figure 2(b) and previous findings (Fortunato et al., 2008). We also introduce a random crawling baseline, where the scorer assigns random scores. We run both of them in a crawl-then-select setting, first crawling 1× or 2× documents and then selecting the top 1× (20M) documents based on scores assigned by the DCLM fastText classifier or the FineWeb-Edu classifier. This process mimics existing data-filtering pipelines, which begin with crawled documents and then apply filtering (Li et al., 2024; Penedo et al., 2024).
@@
We introduce an oracle selection run in which we directly apply the pretraining influence scorer to the ClueWeb22-A dataset and randomly sample 20M documents which are scored as the 10% for pretraining, serving as the upper bound.
@@ 4 Evaluation Results
In this section, we first present the overall performance of Craw4LLM (Section 4.1), followed by further analysis (Section 4.2).
@@ 4.1 Overall Performance
In this experiment, we compare the performance of Craw4LLM with baseline crawlers by evaluating LLMs trained on their respective crawled data. As shown in Table 1, when all methods crawl the same amount of training data (1×), Craw4LLM significantly outperforms random crawling and indegree crawling using either DCLM fastText or FineWeb-Edu classifier as the pretraining influence scorer. In the crawl-then-select setting, where traditional crawlers are allowed to collect twice as much data (2×) for later selection, they still underperform compared to Craw4LLM. This suggests that incorporating pretraining-oriented signals early in the crawling process is more beneficial than relying on post-selection. With only 1× of the data, Craw4LLM retains 95.3% and 95.6% of the performance achieved by the DCLM fastText and FineWeb-Edu oracle run, respectively, which directly selects from the entire ClueWeb22-A dataset, a substantially larger 45× data pool.
@@ 4.2 Analysis
In this subsection, we further analyze the efficiency of Craw4LLM compared to traditional crawlers and explore the reasons behind it. We utilize the DCLM fastText classifier in the experiments presented in this subsection.
@@
We evaluate the efficiency of Craw4LLM by comparing the number of documents it crawls or visits against baseline crawlers. As shown in Figure 3(a), even when the baselines crawl 4× the required pretraining data for selection, they still underperform compared to Craw4LLM. Extrapolation suggests that the indegree-based crawler would need to crawl 4.8× documents (96M) to match Craw4LLM’s performance. Figure 3(b) further illustrates that Craw4LLM achieves the same performance while crawling only 21% of the documents required by the indegree-based crawler, or 48% when considering all visited documents. These results highlight the efficiency of Craw4LLM, demonstrating its potential to reduce website burdens and mitigate over-crawling.
@@ Appendix A Details on Crawling
We run our simulated crawlers on a Linux server equipped with two Intel(R) Xeon(R) E5-2630 v3 CPUs (8 cores per socket, 16 cores in total, 1 thread per core), 125GiB of memory, and an SSD. The crawling process of Craw4LLM takes about 21 hours to finish. In comparison, the random and indegree-based crawlers take around 10.5 and 12.5 hours, respectively. Note that since these are simulated crawls on the snapshot, the reported times do not reflect real-world crawling performance.
@@ Appendix B Details on LLM Training and Evaluation
Appendix C Details on the Pretraining Influence Scorers
DCLM fastText Classifier
(Li et al., 2024) is a quality filter based on fastText (Joulin et al., 2017), trained to distinguish between high- and low-quality pretraining data. A dataset size of 400K examples (200K positive, 200K negative) is used for training. Negative samples are random documents from an earlier version of the RefinedWeb reproduction from the DCLM team. Positive samples come from OpenHermes 2.5 (Teknium, 2023) and high-scoring posts from the r/ExplainLikeImFive (ELI5) subreddit. The final DCLM-baseline dataset is created by selecting only the top 10% of documents ranked by the classifier.
FineWeb-Edu Classifier
(Penedo et al., 2024) is a model designed to identify and filter educational content within FineWeb. Based on the Snowflake-arctic-embed-m [^4] https://huggingface.co/Snowflake/snowflake-arctic-embed-m text embedding model, it is trained on 450,000 annotations generated by Llama-3-70B-Instruct, which assigned an educational quality score from 0 to 5 to every document. FineWeb-Edu is built by filtering out samples from FineWeb with scores lower than 3 with the trained classifier.
Appendix D Detailed Results
Pretraining Selection Commonsense Reasoning
Method Influence Scorer Pool Size CommonsenseQA COPA OpenBookQA PIQA
Oracle DCLM fastText 45× 0.2850 0.7000 0.3300 0.6812
Oracle FineWeb-Edu 45× 0.2342 0.6200 0.3620 0.6638
@@ Appendix D Detailed Results
Random DCLM fastText 0.2588 0.6200 0.3160 0.6785
Random FineWeb-Edu 0.2301 0.6200 0.3140 0.6779
Random DCLM fastText 0.2326 0.6400 0.3380 0.6757
@@
Indegree DCLM fastText 0.1966 0.6600 0.3040 0.6752
Indegree FineWeb-Edu 0.1974 0.6400 0.3160 0.6616
Indegree DCLM fastText 0.2088 0.6400 0.3400 0.6817
Craw4LLM DCLM fastText 0.2277 0.6600 0.3300 0.6926
Craw4LLM FineWeb-Edu 0.3219 0.6200 0.3500 0.6616
@@
Pretraining Selection Language Understanding
Method Influence Scorer Pool Size BIG-Bench Lang. Id. HellaSwag (zero-shot) HellaSwag LAMBADA Winograd Winogrande
Oracle DCLM fastText 45× 0.2515 0.3856 0.3905 0.4432 0.6557 0.5130
Oracle FineWeb-Edu 45× 0.2580 0.3831 0.3818 0.3643 0.6227 0.5185
@@
Random DCLM fastText 0.2468 0.3882 0.3925 0.4073 0.6007 0.5130
Random FineWeb-Edu 0.2485 0.3815 0.3800 0.3804 0.6007 0.5146
Random DCLM fastText 0.2521 0.4011 0.4019 0.4390 0.6154 0.5130
@@
Indegree DCLM fastText 0.2547 0.3749 0.3771 0.3773 0.5861 0.5241
Indegree FineWeb-Edu 0.2528 0.3636 0.3658 0.3672 0.5678 0.5193
Indegree DCLM fastText 0.2562 0.3994 0.4008 0.4159 0.6190 0.5178
Craw4LLM DCLM fastText 0.2544 0.4035 0.4048 0.4196 0.6593 0.5288
Craw4LLM FineWeb-Edu 0.2521 0.3726 0.3717 0.3478 0.6264 0.5083
@@
Pretraining Selection Reading Comprehension
Method Influence Scorer Pool Size BoolQ CoQA SQuAD
Oracle DCLM fastText 45× 0.5755 0.2479 0.3139
Oracle FineWeb-Edu 45× 0.5367 0.2305 0.3130
@@
Random DCLM fastText 0.5807 0.2053 0.2759
Random FineWeb-Edu 0.5627 0.1997 0.2304
Random DCLM fastText 0.5911 0.2361 0.2951
@@
Indegree DCLM fastText 0.5697 0.1843 0.2390
Indegree FineWeb-Edu 0.5670 0.1904 0.2011
Indegree DCLM fastText 0.5765 0.2147 0.2736
Craw4LLM DCLM fastText 0.5440 0.2264 0.2215
Craw4LLM FineWeb-Edu 0.4654 0.2146 0.3026
@@
Pretraining Selection Symbolic Problem Solving
Method Influence Scorer Pool Size AGI Eval LSAT-AR BIG-Bench CS Algorithms BIG-Bench Dyck Lang. BIG-Bench Operators BIG-Bench Repeat Copy Logic
Oracle DCLM fastText 45× 0.2739 0.4341 0.2160 0.2143 0.0625
Oracle FineWeb-Edu 45× 0.2826 0.4606 0.2870 0.1762 0.0313
@@
Random DCLM fastText 0.2696 0.4538 0.2520 0.1762 0.0313
Random FineWeb-Edu 0.2000 0.4091 0.2050 0.1476 0.0313
Random DCLM fastText 0.1957 0.4568 0.2600 0.1857 0.0625
@@
Indegree DCLM fastText 0.2609 0.4235 0.2340 0.2143 0.0313
Indegree FineWeb-Edu 0.2304 0.4545 0.2060 0.1619 0.0313
Indegree DCLM fastText 0.2174 0.4538 0.2530 0.1667 0.0938
Craw4LLM DCLM fastText 0.2696 0.4371 0.1620 0.2095 0.0938
Craw4LLM FineWeb-Edu 0.2609 0.4750 0.1780 0.2048 0.0938
@@
Pretraining Selection World Knowledge
Method Influence Scorer Pool Size ARC Easy ARC Challenge BIG-Bench QA Wikidata Jeopardy MMLU
Oracle DCLM fastText 45× 0.5951 0.3166 0.4945 0.1176 0.2805
Oracle FineWeb-Edu 45× 0.6343 0.3549 0.5407 0.1726 0.2706
@@
Random DCLM fastText 0.5425 0.2807 0.5081 0.0648 0.2561
Random FineWeb-Edu 0.5535 0.2816 0.5194 0.0863 0.2483
Random DCLM fastText 0.5577 0.2867 0.5126 0.0970 0.2543
@@
Indegree DCLM fastText 0.5248 0.2790 0.5205 0.0555 0.2464
Indegree FineWeb-Edu 0.5143 0.2679 0.5044 0.0553 0.2389
Indegree DCLM fastText 0.5749 0.2935 0.5084 0.0959 0.2430
Craw4LLM DCLM fastText 0.6103 0.3208 0.5143 0.1323 0.2661
Craw4LLM FineWeb-Edu 0.6427 0.3592 0.5319 0.1859 0.2736
@@
Appendix E The ClueWeb22 Dataset